Standard Rules

for MIT Assassins’ Guild games

2004–2005 edition

\begin{matrix}c \\ ◯ \end{matrix}2004 Jamie Morris (editor) and the MIT Assassins’ Guild

Preface This document is presented as a set of Standard Rules for Assassin games; that is, a compilation of the most-used rules of actual Guild games. No claim is made that these rules are best in any way nor that they have prescriptive force—just that they are, in practice, standard. Most of them are so standard the SR presents them just as rules; some say “By default… ” or “Unless otherwise specified… ” because while GMs will often want to leave it at that, they’ll often want to do something more complicated.

Revisions occur each summer to reflect what games have been doing; yearly changes are expected to be small—see Appendix B Edition History. They are discussed on the rules-discussion mailing list first. Send comments and suggestions there or to rules-editors@mit.edu.

   The SR live at http://web.mit.edu/jemorris/Rules/ (on Athena, /mit/jemorris/Rules/).

Playing with the Standard Rules If you’re playing in a game with the Standard Rules, that means the GMs have declared this the basis of their rule set. Since players are always responsible for knowing the rules, you’ll need to know these. Keep in mind that the GMs’ game-specific rules document takes precedence if there’s any conflict. (It should be clear from the game rules when this happens.)

Since changes only happen annually and tend to be small, you won’t need to reread the SR entirely for every game. But please do for at least the first few, not just once! It’s roughly ordered by complexity.

If you’re already familiar with a previous edition, the parts you’ll need to check are marked with changebars (and summarized in Appendix B Edition History); occasional rereading should suffice for the rest.

GMing with the Standard Rules Games of any length and size can use the SR. If you choose to, just have your game’s rules proclaim the Standard Rules to be in effect, with a pointer, and drawing attention to which of your rules (if any) override SR stuff. A template is available with the SR.

It is suggested you divide your rules seminar into an early part for less experienced players, at which you go over the SR (hopefully having copies for them), and an all-players part for your game-specific rules.

Assassin Games The Game Masters (GMs) wrote and are running the game—they’re in charge. When you have any doubts about the rules or how the GMs interpret them, talk it over with a GM in advance. If you make a mistake because you didn’t read the rules or didn’t ask a GM to clarify, you’re out of luck. When you’re unsure whether something is legal and haven’t asked a GM, assume it’s illegal and don’t do it.

Please note that many of the rules are almost impossible to enforce, so the game relies heavily upon the honor system. Remember, it’s only a game. It’s supposed to be fun. We hope that you agree that it is no fun to win by cheating. Be your own harshest critic.

Playing loopholes in the rules is also frowned upon. This is a role-playing game, not a law school. The GMs reserve the right to make decisions which violate the letter of the rules in order to preserve their spirit.

All GM rulings are final and override the rules. You may disagree with a ruling, you may not like the rules, but they apply equally to everyone. Remember the GMs may be considering more than you know.

While the GMs are friendly folks, the sometimes competitive nature of the game will occasionally cause them to be stern as well. Don’t take it personally. They’re trying to keep game running smoothly and fairly.

The GMs are also available to answer any questions about your Character, how the game works, whatever. They know your secret plots already, so you can talk freely; use them as a resource, it’s one of their jobs.

Remember, the GMs are human. They’re also probably very stressed by the effort of writing and running this game! Be nice to them. Yes, they can make mistakes, and they can make rulings inconsistent with game sheets or each other. They might have a good reason, or they might have just screwed up. It happens; you should either live with mistakes or point them out very politely, very calmly, once. Then keep on playing.

Assassin games are often tightly plotted. People may depend on your presence and actions in order to complete their own plots. If you are not going to spend any time in game, please turn in your Character Packet so the GMs can give it to someone else. If you are unexpectedly hosed, please try to warn the GMs and your allies.

1 Reality and Game Reality 1 1.1 Non-Players . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Non-Player Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Observers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.4 Mechanics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.4.1 Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Character Packets 2

3 Game Times 3 3.1 Halts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2 Not Here . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 Game Space 4 4.1 Restricted Areas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.2 Living Areas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4.2.1 Jhereg Rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5 Items 5 5.1 Bulkiness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.2 Bodies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.3 Containers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5.4 Ingestibles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5.5 Written Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5.6 Finders Keepers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5.7 Searching People . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

6 Violence 7 6.1 Health . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.2 Zone of Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.3 Martial Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

6.3.1 Using Weapons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.3.2 Fleeing Fights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

6.4 Ranged Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.5 Stealth Abilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.6 Waylay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.7 Weapons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.8 Explosives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.9 Hapless Victims . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6.9.1 Killing Blow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.9.2 Bonking Heads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.9.3 Captivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.9.4 Interrogation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6.10 Death . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

7 Miscellaneous 10 7.1 Administration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 7.2 Athena and Phones . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

A Combat Systems A.1 A.1 MA / MD / HP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . A.1 A.2 Rock, Paper, Scissors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . A.2 A.3 Phi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . A.3 A.4 Drive-By Darkwater . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . A.4 A.5 Melee Darkwater . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . A.5 A.6 Cinematic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . A.6

B Edition History B.1 1 Reality and Game Reality This is only a game. Remember the difference between reality and game reality. This

game does not give you the right to do anything normally prohibited—hacking is right out.

You yourself are a Player. In the game, you are playing the part of a Character. (Specifically, a Player Character, or “PC,” as opposed to an “NPC” (see 1.2 Non-Player Characters).) To some extent, this will require you to pretend that you have a split personality—your Character may have different tastes, background, knowledge, and morals than you, the Player. Think of yourself as an actor taking on a role.

At times during the game, emotions may run high. You may find that your Character has gotten into a horrible argument with another Character and is quite angry. This is OK; if you’re really role-playing well, this can even be a lot of fun. If you find that the argument has crossed that Player / Character line, and you, the Player, are getting angry with the other Player, this is not acceptable—it is no longer fun.

If you notice yourself crossing this line, take a break for a little while to calm down. If you think someone else is crossing the line, you might politely suggest that they do the same, but don’t push the point. Some particularly good role-players can portray anger VERY convincingly without actually feeling any.

If someone politely suggests that you take a break to calm down, please consider doing so. Perhaps you crossed the line without noticing it.

There is a sharp distinction between Player reality and Character reality. Players are expected to treat each other with courtesy and to explain to other Players what their Characters perceive in confusing situations (“My Character’s hands are covered in blood.”) Characters are under no such restrictions, and should do what it takes to further their goals (“Uh, hi Bob. Just got back from the butcher shop.”) As this example indicates, while Characters may lie to other Characters at almost any time about anything, Players may never lie to anybody about anything.

1.1 Non-Players There are lots of people around who aren’t in the Guild, some of whom actively dislike us. Annoying these people is impolite and can get the Guild in trouble. Ideally we’d avoid them ever noticing that a game is happening; in practice we do what we can to be less intrusive and less disturbing.

Disturbing behaviors include: being noisy or running around, especially near labs and offices; blocking hallways (this only takes a few people); and using threatening body language, having conversations featuring “bomb” or “kill them all,” chasing each other, screaming for help, or otherwise alarming nervous people.

Never have combat around NPs. Avoid having weapons be visible to them. Guns, even plastic ones, sometimes disturb people, and foam boffers can look like lead pipes at a distance. Defeating your enemies is less important than not bothering NPs, even when you were about to spring the perfect ambush.

Since NPs aren’t part of the game, you can’t deliberately use their presence to protect you from attack— rather, you should actively avoid such protection. So while a quiet, non-threatening conversation is in itself OK to have around NPs, it’s still better taken elsewhere since someone coming upon you (or among you!) may wish to do more, like kill you all. Worse: they may not notice the NPs and actually try it.

When someone (especially the Campus Police) asks what you’re doing, tell them you’re playing a game with the MIT Assassins’ Guild. This will disturb them much less than if you’re evasive; after all, you aren’t doing anything wrong. If they persist, direct them to the GMs. Be polite and tactful.

If, despite your most valiant efforts, some NPs do get upset, call the GMs to help calm them down! Complaints to the Dean’s Office are to be avoided at all costs.

You are encouraged to spread the gospel of real-time, real-space role-playing to your friends. However, NPs cannot be used for any ingame effects. They can’t hold items, pass information, or spy for Players, nor can you hide things in their rooms, and they most certainly can’t help you kill someone.

1.2 Non-Player Characters NPCs do exist ingame, but are played (at least temporarily) by a GM or a helper. Unlike NPs, they can be spoken to, bribed, attacked, etc.—they’re much like PCs. Some aren’t readily distinguishable from PCs, while others essentially just provide mechanics. They may or may not be stupid, unable to report events or identify people, impossible to persuade out of their tasks, or otherwise limited.

1.3 Observers An Observer (or “ghost”) is an official spectator. They’re not participating, but the game goes on

normally in their presence. (If you’re unsure whether someone is an Observer or an NP, assume NP.)

Observers are usually required to wear white headbands (and maybe “Observer” badges) in game areas. Anyone who wants to observe should contact the GMs to get one and learn what other rules may apply.

The GMs may use some Observers to do helpful things such as carry messages or implement mechanics. When they’re just hanging around, they shouldn’t get in your way, and must leave if you ask them to.

1.4 Mechanics Many Character actions are represented by the same Player action: for your Character to walk and talk your Player just walks and talks. Other things, such as torture or brainwashing, must be simulated. A mechanic defines how some aspect of game reality is implemented.

1.4.1 Numbers Often mechanics information is encoded in numbers on things like character badges or item cards. If the GMs tell you how to decode them, what you learn is ingame knowledge, but the mechanics and numbers themselves are not. You cannot discuss numbers or encodings (even public ones), keep lists of numbers, etc. (Say “she looks elvish” or “this painting is a masterpiece” but not “since her badge number has a 7” nor “since its item number is prime.”) Do not try to figure out mechanics the GMs have not given you.

2 Character Packets Your Character Packet typically comes in a big manila envelope. It has your role—who you

are, what you’re up to… everything about your part in the game. Read all the contents thoroughly.

For purposes of game, Character Packets do not exist and are confidential. Therefore, none of the contents, unless clearly stated otherwise (the main example being the cards representing items), should ever be shown to any other Player under any circumstances. This also means that you should never look at anything from someone else’s packet. It’s usually best not to keep game items in Character Packets, since people should be able to find your items but shouldn’t search your packet.

Even if you’re positive someone else has an identical sheet, you still can’t look at each other’s, nor should you talk about “well, our Greensheet says… ”—your sheets don’t exist ingame, so Characters shouldn’t discuss them, and you might be wrong about having the same one (GMs love to play tricks).

Some things often found in packets are:

Badge Always wear yours visibly when in game areas. Use the normal-printing side when you are present in-Character (“Here”). It may have your Character’s name, description, etc, and/or your Player’s name. Character names and descriptions are ingame information; badge numbers are not (see 1.4.1 Numbers). The reverse-video “I’M NOT HERE” side is for when your Character isn’t with you (see 3.2 Not Here).

Your badge holder is often a convenient place for out-of-game things like stat cards, ability cards, and memory packets. It’s not a good place for game items.

Character Sheet This has your personality, history, beliefs, goals, etc—who you are, what you’re doing, why you’re doing it, who your friends and enemies are. … It is the single most important thing in your packet. Usually your other sheets add more information, but your character sheet is always dominant. Become very familiar with it. Never show this to another Player.

Your charsheet probably also has a list of what else your packet contains. See the GMs if you seem to be missing anything or to have things you shouldn’t.

Bluesheets These are like character sheets for an entire group of people. They contain goals, motives, history, etc., common to a group. What this actually tells you is what you think group members think. Bluesheets do supplement your Character sheet, but Character sheets always take precedence. Some group members may have radically different views, maybe even you. Never show Bluesheets to another Player.

Greensheets These give you special knowledge and/or tell you how some of your mechanics work. If you’re not sure how to use it or what it’s good for, talk to a GM. There may be some public Greensheets. Never show a non-public Greensheet to another Player.

Statistics Card This has various information about you, much of which is meant for mechanics, often ones you don’t have, so you often won’t know what it means. Keep your stat card with you in case you encounter such mechanics, but never show it to another Player.

Any real-world statistics (like age or nationality) are known to you ingame and can be discussed freely. Stats for public mechanics (or private mechanics you happen to know about) can only be discussed directly if the GMs say so—otherwise, you can only say “I’m really pumped!” or “My telepathy is fairly weak” but not “I have a 17 combat score” nor “My μ is 23.” Whatever discussion is allowed need not be honest. …

Any other statistics are complete mysteries. Since your Character doesn’t even know you have an alpha stat, you can’t even speculate about what (if anything) it represents, much less discuss it ingame.

The back of your stat card may have a Death Report Form (see 6.10 Death).

Ability Cards These represent special abilities your Character has. One side tells you how the ability works and is never shown to anyone else. The other side is shown only to those you use it on, telling them what to do and/or what their Characters experience: “Tell me your phi stat.” “You are paralyzed for ten minutes.” “I have seduced you. Come have sex with me for fifteen minutes; we may not hurt each other for an hour.” Let bystanders know what their Characters see, if anything. Not all abilities are used on others.

Check with the GMs if you aren’t sure about what your abilities can do. If you have any you might want to use in combat or other time-critical situations, become familiar with their specifics before game starts! Combat won’t stop for people to figure out their character packets.

Memory Packets These are usually small manila envelopes or folded-and-stapled strips of paper. When you encounter the trigger written on the packet, open the packet and read the contents. These may be a suddenly-recalled memory, a flash of ESP, being mind-controlled, whatever; your Character may gain information, react compulsively, and/or be otherwise affected. Triggers can only go off during game times.

Explicit triggers saying “open if… ” go off as specified.

Numerical triggers go off upon seeing that badge number, item number, etc. (i.e. an out-of-game number). Phrase triggers go off when your Character hears that said or sees that in writing (i.e. in-game words). Odd triggers like “Q” go off when something specifically says so (i.e. a sign, ability card, or other mechanic

instructs the Player “open your Q packet now”).

Characters don’t know about memory packets, so you can’t let having (or lacking) unopened ones affect your game actions. In particular, you can’t seek out or avoid triggers, nor maneuver people into saying them. Memory packets can’t be shown to anyone, stolen, etc, whether opened or not.

Once opened, they’re like your Character sheet: any information provided is ingame, so you can tell people “I just realized… ” (though never “my memory packet says… ”). You can generally assume your Character is also aware that the realization or impulse was inspired by whatever you saw or heard.

If a memory packet has an effect on you (“fall asleep,” “lose two μ,” “say ‘hovercraft’,” “obey this man”), assume it functions every time you encounter the trigger, unless told otherwise. If in doubt, ask a GM.

You cannot call a halt or otherwise interrupt game for memory packets; you can be shot while reading. Usually when a memory packet is triggered you can wait to finish your conversation and go around a corner before reading it, though you should open it immediately if the trigger is an event or time (or says to). If you see someone else reading a memory packet, your Character notices their Character looking very thoughtful.

Items These do exist ingame and can be shown to others, traded, stolen, etc. See 5 Items.

Whitesheets These are ingame written documents—blueprints, contracts, newspapers, maps, rituals. … The Whitesheet has the actual information: what the Player sees includes everything the Character sees, whereas an item card for a book may resort to “a history book, see a GM to read” since cards are small. Whitesheets should always be marked as game items. As such they can be shown to others, copied, counterfeited, traded, stolen, destroyed, etc. See 5.5 Written Information.

Playerlist Just what it says. Please respect Players’ “don’t call” times. Changes to the playerlist are likely; the GMs will propagate the new information. If you’re listed incorrectly, tell the GMs.

Scenario This provides background information like in-world history, recent events, and important rooms. It sets the scene and perhaps mood for the game. Keep in mind that a Scenario describes public knowledge and perceptions, which are often incomplete or outright wrong.

3 Game Times The GMs will specify when game begins and ends. No game action may happen before game start— no talking, no interacting, no searching areas, no hiding items. It is highly recommended that you take the time to become thoroughly familiar with your character packet.

Long games often have a Game Break in the middle. During Game Break no game action may take place, including in-Character communication. Catch up on your tooling.

Game Breaks typically run for about 36 hours, like 6am Tues to 6pm Weds. Unless stated otherwise, assume that one game day passes during this time for purposes of healing, regaining magical energy, etc, though addictions and so on do not worsen during this time.

3.1 Halts Game Halts are for when you need to resolve Player confusion, want to get a GM on the scene immediately, or spot an NP coming through (see 1.1 Non-Players). They’re called by saying “Game Halt” loudly enough for all Players in the area to hear.

During a Halt NO game action may occur. You may not talk about any aspect of the game, except in regards to resolving the current situation. People should not move around or change things; the situation when the game resumes should be the same as when it left off. If other Players wander in, they become Halted as well; a GM may well be needed to sort out the multiple Halt. This is one of the reasons to keep Halts to a minimum. Summon your GM via Observer when possible.

A Halt is terminated by consensus of all Players involved in the Halt, or by a GM. Someone should call “3, 2, 1, Resume” so it is clear exactly when play resumes.

Halts may not be used to stall, gain thinking time, etc. Use sparingly. Use common sense. Hopefully, if any Halts are necessary, they’ll be resolved very quickly. Be reasonable and be familiar with the rules.

3.2 Not Here In some instances, your Player may need to do something or go somewhere the Character isn’t; in that case, you can go Not Here by flipping your badge to the Not Here side. Tell other Players that you see that you’re Not Here, so they know you’re not currently playing. Your Character does not see anything your Player sees while Not Here; therefore, it’s best if you just avoid game area entirely when Not Here. Not Here is a good way to do homework, or to run for a GM when you’re in the middle of a game-halt; you should NOT use it to leave game in the evening without first walking to a public place (street, lobby 7, …) because someone might be trying to follow you and they deserve the chance.

Times to go Not Here are: when you are out of the playing area; when your Character cannot be there (he’s off being brainwashed, for example); when you are doing real-world work and shouldn’t be disturbed; when an ability, assistant, or GM tells you to. See 6.10 Death for an important case.

It is important to understand Not Here and not to abuse it. Ask yourself, “Do I need to be Not Here for this?” If the answer isn’t Yes, and if you’re not going to fetch a GM, then be Here. If the Player is going to the bathroom, so is the Character… and someone can follow and shoot you in the hall.

When you are Not Here, you may not act on anything you see or interact with the other Characters in the game. You cannot become Here based on anything you see or anyone you meet, although you can become Here based on, say, phone calls or messages (see the section on Athena). You do not need to be Not Here to go about your normal day-to-day activities unless you want to. If someone asks “Are you Here?” it is never ok to say “I can be”; you are or you aren’t—no giving away the information that you’re alive and free without also being Here to be shot.

4 Game Space Game action may take place only in game areas. This covers anything done by Characters, including ingame discussions and stashing items. The only exception is that phone and computer use remain as permissible outside game space as they ever are (see 7.2 Athena and Phones). When in a non-game area, you are automatically Not Here (see above).

Bathrooms, activity offices (SIPB, LSC, MITSFS, etc), private offices, and places that are sometimes locked (except living areas) are never ingame. Anywhere that Players are not allowed to be (e.g. roofs, steam tunnels) is not ingame. All MIT students must have constant legitimate access to game locations.

Athena clusters are only ingame for purposes of talking and using Athena, and those only if the game allows Athena use at all. No other action may occur there. Try to keep the talking down, as that’s not what clusters are for. Do not use these rules as a shield to give you a safe place to hide or communicate.

Anywhere off-campus is not ingame, except Player living areas (for which see below).

4.1 Restricted Areas Many games will declare parts of campus to be ingame but not normally accessible. This

titled

means you can’t just enter or leave there freely, but you might be able to under some circumstances—it’ll be made clear if it happens—at which point game action can take place there normally.

For instance, a game might restrict everything above the third floor and have those areas represent secret catacombs whose concealed entrances you must discover. These may not be contiguous, and may further subdivide and/or interconnect in ways unrelated to the real-life buildings.

Should you find that you can go to one of these areas, follow the directions given for doing so quickly and efficiently. You are Not Here while in transit to and fro, so may take no game action, except that you may still speak with your companions.

If your Player must enter such a restricted area for Player reasons (e.g. to get to a class or an office) you must be Not Here, so try to avoid game-related signs etc since your Character doesn’t see them.

4.2 Living Areas When game area is basically campus, Player residences are by default also Character residences

and hence legal locations. Game action is still subject to the NP rules and common courtesy.

Some Players live off-campus at ILGs or real-world places. These are all considered to be on the fringes of campus (game area) for game purposes; the interior of TEP exists, but the distance between it and MIT does not. If you want to take some action at an off-campus location, and you do not have the time or the means to get there (some Players live in Somerville), find a GM and they will try to find a way to kludge it.

Note that if you hide items in your room, further concealment is pointless. Since Players searching your room will not be physically searching your room, it is automatically assumed that anyone who manages to search another person’s room finds all the items there, unless special abilities come into play.

4.2.1 Jhereg Rule Any Player may retreat to his or her room to study, sleep, or whatever in safety.

The Jhereg Rule (named for Steven Brust’s series) is this: a Character may not be present ingame in another Character’s room unless invited there ingame. That is: if you pass by a Player’s room and see the door open and the Player nowhere around, you may not physically enter and search. If you go over to another Player’s room and they open the door for you you may not enter without permission. Furthermore, regardless of whether you the Player are invited into the room, you may not take any game action until invited to do so by the other Player. You may not, standing in or out of the room, shoot someone who is in their room, unless the other person has started discussing game business with you, performed game actions on you, etc, in which case they forfeit this protection. Do not abuse this; your Characters are unaware of it. It gives you safety to go over to your friend’s room to work on math homework without either of you getting shot. If the Player whose room it is starts discussing game all bets are off. Shoot away if you like. Visitors may ask to talk to the room owner’s Character, but the room owner may decline.

If you want to go to someone’s room specifically in-game, you should indicate this by something like using your Character name when they ask who’s there. If they open the door for your Character you are considered invited to conduct game business.

The Jhereg rule is to allow Players to have lives outside game. It cannot be used as a shield to take game actions behind the invulnerable barrier of your door, especially with other Characters. If game action is occurring inside, PCs outside can get in however the GMs feel appropriate (e.g. busting down the ingame door); it’s best to just avoid doing things people would want to break in on in your room.

Rooms with multiple roommates playing are defined to be in different rooms for this mechanic.

5 Items Items are mechaniced as little white cards marked “Game Item.” They can be exhibited, traded, stolen, etc.

Some items have associated props—physreps, Physical Representations. Always keep these together with their item cards (preferably taped on). If they’re separated, the item card counts as the actual item.

Guns and ammunition are the only items typically without associated item cards. Guns often have item stickers or small slips of paper instead; stick these to the physrep for it to represent an ingame gun. Ammunition is by default trivial to get and has no item card; if you have a gun, you can just load it.

The description on an item card (“a large, glowing, blood-red ruby”) is ingame information obvious to anyone who looks at the item. Item numbers are not ingame (see 1.4.1 Numbers) and may not be unique.

There may be explicit restrictions on carrying some items (see below). Even when there aren’t, be reasonable with items. You can’t carry a hundred rocks in your pocket, you can’t fold a sword in half, you can’t hide a life-sized statue in a fire hose. Your Character shouldn’t either. Only do things with item cards that your Character would be able to do with the actual items; use common sense.

If you wear an item (e.g. jewelry), display the item card (perhaps taped to badge holder) as well as any physrep—someone might have a memory packet or other mechanic requiring the item number or description.

An item may have an envelope attached (or simply be an envelope with item number and description). The most common cases are in 5.3 Containers and 5.4 Ingestibles. In other cases, it will have either a code or instructions. If a code, ignore the envelope unless you know otherwise (e.g. you have a memory packet or special ability for it). Any instructions should be self-explanatory (“open if you push the button”).

5.1 Bulkiness A bulky item is too big or heavy to be carried freely. Bulkiness is measured in “hands,” i.e. how many hands it takes to carry it. An item card with N colored dots on it is N-hand bulky. (A dot with a number on it is that many hands’ worth of bulkiness.)

Cooperation is possible. It should be clear to onlookers which people are carrying the item (the Players should be holding the item card together).

A bulky object must be carried openly and cannot be concealed, except when in a container (below) with enough capacity. Hands helping to carry a bulky object can be doing nothing else.

N − 1 hands may drag an N-hand bulky object at a slow pace.

5.2 Bodies Characters are two-hand bulky items, with badges instead of item cards and Players as physreps.

To move an immobilized, unconscious, or dead Character, you must carry or drag them. As with moving any other bulky object, it should be clear to observing Players what is going on. Typically a single carrier will walk behind their burden with hands on their shoulders, or two carriers will walk to either side with hand on shoulder. If the Player isn’t there, carry their badge conspicuously. Tell any Players who see you what their Characters see. (“We are carrying an unconscious body”—they can’t tell if it’s dead without close examination, unless it’s headless or some such.) You do not, of course, ever carry a Player.

5.3 Containers Crates, bags, and other container items have an envelope or other physrep for holding their contents and are marked with a capacity specifying how many hands of bulkiness fit inside. The bulkiness of the contents is effectively negated; only the container’s own bulkiness is relevant to moving it. As with pockets, more non-bulky items can always be stuffed in, but be reasonable.

Player bookbags, purses, etc are only non-zero-capacity containers if they’ve got an item card saying so.

5.4 Ingestibles Anything that you can eat, drink, or be injected with in game will be mechaniced by a small packet saying, for instance, “A syringe filled with clear liquid,” or attached to an item card saying that. If you eat, drink, or otherwise take the stuff into your body, read the contents. Once you actually open the packet you are committed, so think carefully.

If you take two ingestibles with similar effects, assume you are affected only by the stronger unless told otherwise. If eating a piece of cake makes you feel good, eating two doesn’t make you feel twice as good. (Alcohol is a common exception.) You experience simultaneous effects from ingestibles that don’t overlap.

If you take a drug with the same effect before the first has worn off, reset the clock on its effects.

5.5 Written Information All paper should indicate whether it is a game item / ingame or not.

White paper (playerlists, maps, newspapers, Whitesheets, etc) is almost always ingame; rules are not. You can write “gamename Game Item” on unused white paper to make it ingame. Like other game items, ingame paper can be traded, stolen, lost, etc, so don’t write your secrets on it!

Colored paper (like Bluesheets) is almost always out-of-game. Charsheets, Greensheets, memory packets, etc, are always out-of-game regardless of the paper. You can write “gamename NOT IN GAME” on unused paper, preferably colored. Don’t keep this in a pocket with game items! (Try your badge holder.)

Information gained by your Character ingame (or on your packet’s pregame items) may only be written on ingame paper. Only write on your out-of-game papers information from your other out-of-game papers, from GM rulings, and from publicly-GM-supplied material such as a playerlist.

5.6 Finders Keepers Items can be carried around or hidden. To hide things, you must actually hide the item card (and any physrep) somewhere. If someone finds and takes it, oops, try to figure out how to get it back. … You can only hide things in game areas. Be sure to put bulky items somewhere appropriately spacious. Often you can leave stuff in your room; see 4.2 Living Areas.

If you find game items someone else left lying around, by all means take them, they’re yours now.

To steal items left in someone’s bookbag or purse, do not search the Player’s belongings. Leave a “STOLEN: see a GM” note on it and tell a GM. The Player will give any ingame contents to the GMs, who will pass them on to you.

To steal game items left elsewhere, you must actually search the area. If you find something, wonderful; if you don’t, oh well. Don’t bother trying to search non-game locations; even if it weren’t illegal, it wouldn’t do you any good since no one’s allowed to hide things there.

To search a game location your Player does not have access to, such as a person’s room, find a GM and convince them that your Character can get in. This may or may not be possible without special skills (like Lockpicking). If you’re allowed to just “break the door down” or “shoot the lock off” and so on, you’ll usually be required to go knock on the door—see 4.2.1 Jhereg Rule for what happens if the Player is in—stand in the hall for a bit so you can get caught, and leave a note like “lock shot off, items stolen, see a GM.” The GMs will pass the items on to you.

5.7 Searching People All searches of Characters (or their belongings) are conducted via Player dialogue.

Only Characters who are willing or helpless can be searched. Even then, anyone nearby can trivially prevent it—see 6.2 Zone of Control.

There are two kinds of searches. If the searcher does not indicate which, they are performing a Total Search by default. It takes two free hands to perform either kind at a reasonable speed.

Pat-Down Searches simulate stereotypical TV cop “frisking.” This search will only reveal the presence of weapons (ranged and hand-to-hand) and armor. This takes as much time as it takes your victim to tell you what you find. (“You find a crossbow in this pocket and a knife in this one.”)

Total Searches simulate a mildly invasive complete search of the Character’s entire body and clothing. This takes as much time as it takes the victim to hand over their item cards. Victims, don’t stall, but do be careful to hand them only item cards and not ability cards and so on. Anyone nearby can see what the searcher finds, which is another reason not to hand over an indistinguishable wad of cards, as the searcher doesn’t find a wad of items. Remember to hand over your ingame money.

Searching a Character does not mean you’ve searched their backpack, purse, etc. For those, tell the Player “I’m searching your backpack,” and unless you’re prevented they’ll give you any game items in it.

6 Violence Combat is strictly ingame. If there’s any risk of real-world injury or property damage, call a Halt.

6.1 Health Normally your Character is fine: conscious, mobile, etc.

new

If knocked out, you’re unconscious but will awaken five minutes after whatever last knocked you out. If wounded, you’re unconscious, bleeding, and will die five minutes after whatever first wounded you. Upon losing consciousness, drop anything you’re holding, fall down, and just lie there. Try to look

obviously unconscious to other Players, and ignore conversation and other game action around you. Helping the injured usually requires a special ability; assume you can’t unless told otherwise. If immobilized, you are helpless and may do nothing but talk.

6.2 Zone of Control Someone is in your Zone of Control (ZoC), and you are in theirs, if you could touch fingertips

with arms extended. ZoC defines Characters being “close enough” for mechanics such as martial combat.

Activities that require a willing or helpless victim—searches, Killing Blow, etc—can also be thwarted by anyone in ZoC of victim or assailant. The rescuing Player just says “I stop you” or “No!” or anything indicating opposition. (“Martial Attack” also works.)

6.3 Martial Combat To attack someone, walk up into their Zone of Control and clearly declare “Martial Attack.” If they do not reply with “Martial Defense” (or “Martial Attack”) within two seconds, they are Surprised and lose the first round of combat. If they do respond, normal combat ensues.

Anyone who says “Martial Attack” to the target after the first attacker, within the two seconds of reaction time and before “Martial Defense” is declared, is helping to attack the target that round. Anyone declaring it afterwards is too late and will only join in the next round. Similarly, anyone in the combat can only declare “Martial Attack” on someone uninvolved, bringing them into the combat, between rounds.

If someone declares Martial Attack on you from outside ZoC, the proper response is to declare Martial Defense and then point out that they were too far away. They may have a pertinent special ability that allows them to attack you from there. Don’t do this unless you know you can.

Martial combat takes place in real time for anyone not directly involved in the combat. Another Character may freely join into the combat on either side (for the next round), fire a gun into the combat, shout for help, or anything else he may choose to do, regardless of how long the combatants are fiddling. Please don’t abuse this; stalling to give your friend time to show up is abusing the rules.

How hand-to-hand combat is resolved varies with the game; it’ll say what system it uses. See Appendix A Combat Systems for the standard ones. General points:

When engaged in martial combat, you may not move around except as required by the combat system. You cannot, for example, walk in one direction pulling the fight towards your friends in the next building.

In multi-person combat, you should point to the person you’re attacking that round. (You almost always get to attack only one person, though you defend against everyone attacking you.) Ganging up is OK, and if you have maneuver choices (e.g. cards, symbols) usually you may coordinate with others (see the system).

Special abilities that work outside combat (e.g. Seduction) almost never work during combat.

6.3.1 Using Weapons Using a weapon requires both its item card and a physrep. Further, the physrep must be in your hand and clear when “Martial Attack” is declared. “Clear” means it’s not in your pocket or tangled up in your coat, even if you’re holding it. It must be free, ready to be used, and visible to onlookers.

You may not draw any weapon after “Martial Attack” is declared (except in the A.6 Cinematic system).

Ranged Weapons If you have your gun already out, shooting is a legitimate response to “Martial Attack.” However, if for whatever reason your shots fail to down them within your two “Don’t be Surprised” seconds, you’re now in martial combat and are Surprised. (You may not both shoot and declare Martial Defense.)

Once engaged in martial combat (having declared Martial Attack or Martial Defense, or having passed the time to say the latter) you may not fire a gun even if you have one drawn.

(Also see 6.4 Ranged Combat below.)

6.3.2 Fleeing Fights To try to escape the combat, say “Break” between rounds. If everyone else says “Break,” the combat ends. If only some do, those who say it are no longer fighting each other but may still be fighting the others. (And next round they can attack each other again.)

If some of your opponents didn’t say “Break” when you did, you’re trying to escape them. You do no damage that round. If you manage to take no damage, you’ve escaped—declare this and leave ZoC.

6.4 Ranged Combat Ranged combat is real-time and based on Player skill at firing and dodging physreps.

totally revised

Never aim at a Player’s head. If your attack hits a Player’s head, your Character hit their Character’s torso. To shoot, stab, or otherwise attack the Character’s head, see 6.9.1 Killing Blow.

Hits on long hair or loose clothing don’t count unless they would have hit the body when passing through. You can’t shield—projectiles intercepted by an item you hold still count as hitting you, not it. The default effect of any kind of projectile hitting you anywhere is to wound you (see 6.1 Health).

new effect ⇒

(When hit location does matter, hits on the Player’s head are still on the Character’s torso.) It’s hard to feel whether and where you’re hit. When there’s any doubt, the shooter calls the shot.

6.5 Stealth Abilities Some Characters can sneak up on people to bonk them over the head, pick their pockets, or

new

otherwise invade their personal space by surprise. Often such things are mechaniced as Stealth abilities.

To use a Stealth ability you have, form the sign of the devil (index and pinky fingers extended, thumb holding other two fingers down) and extend it along a direct, unobstructed line from your shoulder to your victim’s head. Hold this position, staying within ZoC, for the time specified by your ability. You are thwarted if (during this time) your victim notices the symbol or anyone attacks you; otherwise you have succeeded.

Any reaction to the symbol means they’ve noticed; the attacker resolves disputes. When you notice a Stealth attack, make it obvious—“I notice you” is unambiguous. Once the ability’s time is up it’s too late.

6.6 Waylay Waylay is a Stealth ability for mugging people. You do not have this ability unless told otherwise,

new

i.e. by an ability card or by the game rules. It may have prerequisites, such as “must be holding a weapon.” Upon succeeding, say “Martial Waylay.” Your victim is immediately knocked out (see 6.1 Health).

When the martial system includes exotic attack forms, Waylay can probably be used with those as well. See Appendix A.4 Drive-By Darkwater and A.5 Melee Darkwater.

6.7 Weapons The game may make some weapons freely available (often clubs or guns). Unless your weapon is one of those, you must have an item card or sticker (as appropriate) to use it ingame. Further, you always need a physical representation to use any weapon, so that Players can see that you’re a threat to their Characters.

Clubs are physrepped by boffers (foam sticks). “Club” includes anything used as one, like a lead pipe or table leg, so in most situations they’re freely available. If so, their exact effects will be in the game rules, and you’ll be able to use a boffer as a club without an item card (though putting a card on never hurts).

Knives and swords are also foam boffers, swords being longer ones. Tape the item card to the boffer to define the pointy end, so other people can read it when you’re holding your weapon. (If you hold the card end, you’re holding the blade (ouch!) and whacking people with the hilt like a club.) A circle of tape on the blade, ideally brightly colored, will help to distinguish it from a blunt weapon.

Guns, bows, magic wands, etc are usually toy disk guns (which hold twenty shots) or dart guns (which hold one); sometimes more exotic toy guns are used to represent special weapons. These usually have stickers or paper slips instead of item cards, and ammo usually has no item card—see 5 Items.

Throwing knives, darts, and other hand-held missles are usually physrepped by foam pellets, sometimes by ping-pong balls. They usually require a special ability; assume you can’t throw one unless you know otherwise. They have the same effects as other ranged attacks unless otherwise specified.

Weird attacks such as fireballs are most often physrepped as foam pellets. If you see someone holding a foam pellet, you see some indication (their hand glowing, or something) that they may have a weird attack ready. After the weird attack hits, the foam pellet no longer represents anything; you can’t pick the fireball off the floor and throw it back. If you find foam pellets around, give them to GMs.

6.8 Explosives Explosives have two components: the ingame explosive and the trigger.

titled

The ingame explosive is an item card on a foam block, can of play-dough, or similar physrep, visible to Players but not alarming to NPs. The trigger is Player-supplied and should beep or something when it goes off (digital watches are good). Attach it to the ingame explosive. If it fails to go off, so does the bomb.

Generally only people with an appropriate skill can set explosives or disarm them. Unless you know otherwise, assume you can do neither, not even by turning off the real-world timer, nor can you safely move an armed bomb—if you try to throw it, it explodes in your hand.

If you notice such a item start beeping, a bomb just went off. Freeze in position; you are in Game Halt. Figure out what happened, like who’s dead. This may require a GM; if so, hopefully one (or a primed Observer) is already there. Resume when everyone knows what’s happening to them.

Blast radius depends on game and explosive (default to ZoC). Effects also vary; typically anyone nearby (certainly those within ZoC) will suffer the effects of being shot multiple times.

Unless otherwise specified, explosives do not chain-react.

6.9 Hapless Victims

6.9.1 Killing Blow A helpless (or willing) Character may be killed easily by anyone within ZoC by declaring “Killing Blow.” By default, this does not require a weapon. Anyone else within ZoC of you or your target can trivially thwart you as per 6.2 Zone of Control; knocking you down (say by shooting you) also does so. If you aren’t thwarted within two seconds you succeed.

By default you are assumed to have broken your victim’s neck, causing death instantly. Feel free to specify another method, e.g. take a sword and declare “Killing Blow, I cut your head off,” or declare “I’m shooting you in the head” and fire past the Player’s head. Since you never attack a Player’s head, this is the only way to ever shoot/stab/attack a Character’s head.

6.9.2 Bonking Heads Similarly, you may knock out a helpless or willing Character within ZoC (without martial combat) by declaring “I knock you out.” If you aren’t opposed or knocked down within two seconds, your target is knocked out (see 6.1 Health).

k.o.

If a Character is kept unconscious for more than an hour, they may, at the Player’s option, go into a coma or die. It’s boring to be kept out that long, not to mention it’s a real pain having to ignore all the game action around you, like your captors discussing their plans. If you know you’re going to keep your captive out for a while, tell the Player so they can go get a snack in the meantime.

6.9.3 Captivity A helpless or willing Character may be immobilized (see 6.1 Health) by tying them up with

new

rope, duct tape, etc (item cards for which might be freely available, depending on the game). They can get loose if they’re otherwise unimpaired and are left alone for a full five minutes (not poked, nor yelled at, etc).

6.9.4 Interrogation In most games, torture is amusing but inefficient. Unless there’s something special available, you have no way of really forcing someone to give you information. So while you can always roleplay torture, doing so leaves it entirely up to the victim what (if anything) they say and how much truth it contains.

When you’re actually being “truthed”—given Truth Serum, tortured with a special ability, or otherwise having truthful answers forced out of you—or are truthing someone, you’ll want a GM present to resolve ambiguities. Please be reasonable with “loopholes” in truthings; ask the GM before trying to use one.

6.10 Death Death. It happens to all of us sooner or later, and to quite a few Characters during the game. Try to

be a good sport about killing and dying, and you’ll earn the respect of your colleagues.

You should do the following things immediately upon your Character’s death:

• DO NOT give out any information about your Character or their death to any Players.

• Feel free to remain on the scene to play the part of your corpse. It’s nice to do this for at least 10–15 minutes or so, so that PCs who wander along will see you and not have to spot just your badge. Only in this case can you give out any information, and even then, only what’s very obvious. For example, you could tell a Player who looks at you, “I’m lying face-down, not moving, with what looks like a bullet wound in my back.” You can’t say “Joe shot me,” “I was beaten and then killed,” etc.

• When you leave the scene, put any item cards in your possession in a stack where your body is. Put the Here side of your badge on top of it with a description of the obvious state of your body (“sword wound in chest, decapitated”). Take the Not Here side to wear.

• Come tell a GM. If your Stat Card has a Death Report Form, fill that out and give it to the GMs. The form may ask for information like the number of the gun that shot you, so that the GMs can handle forensics; try to include this if you can get it without impeding the flow of play for the survivors.

After your death, you are Not Here (see 3.2 Not Here) if you run into any other PCs. You cannot explain to them why you are Not Here. Dead men tell no tales, nor do Not Here ones. Neither do they talk on the phone, drop hints, or fill their conversations with significant silences. Do your best to communicate no information to anyone (including NPs, as they won’t know not to pass it on to Players). If someone telephones you about game business, tell him the phone is still ringing. If someone sees you in person, tell him you’re Not Here.

If, after your Character’s death becomes generally known to the other Characters, you’re interested in watching the game, come to the GMs and they’ll make you an Observer. In this case you will be given an Observer headband and may freely receive information from Players just like any other Observer, but until the game is over, you may still not convey information to any Player. That means you shouldn’t go asking “why did you kill me?” or other questions that tell them what your Character didn’t know.

Much the same situation applies if your Character is permanently out of game without being dead (e.g. left the space station on a one-way trip, taken to escape-proof jail in another state).

7 Miscellaneous

7.1 Administration The Control Room is where the GMs live, and where you should seek a GM at need. Lots of nifty things you shouldn’t see are there; don’t go in without GM permission. (Knock!) It is a good place to deliver food to.

Often more Players want to see a GM than there are GMs on duty. There will be a Queue near the Control Room, either posted on the Control Room door or written on the chalkboard in the Common Room or Queue Room. The Queue Room is typically the same as the Common Room (below), but may be a nearby room instead; in this case it is probably not a legal game location—that’s why it’s a separate room.

For non-urgent questions, the Box (below) is highly preferable to the Queue.

If you are Not Here and need to see a GM, and there is no out-of-game Queue Room, try to observe the usual Not Here procedures and not see or hear things or talk to Here people.

If you want to just hang out with other Characters, try the Common Room. This is a legal game location. Note that you can be shot there, but that’s the chance you take. The Common Room is generally considered to be NPC-free—that is, no virtual people are there (and no NPs should be in there, either, but treat them normally if they are).

The Box is a rack of file folders that lives in or by the Common Room, Queue Room, and/or Control Room. There will be a folder with each Character’s name as well as “GMs,” and possibly more for various mechanics. Look only in your own folder, unless you know otherwise. If you want to get non-urgent answers from the GMs, leave the questions in the “GM” folder with your Character and Player names on them. The GMs will get back to you, probably by leaving a note in your folder. Make sure you check your folder regularly in case the GMs left something for you.

7.2 Athena and Phones In some games Athena and/or phones don’t make any sense ingame (e.g. medieval fantasies); in any game, they can reduce the live-action component by safely distancing PCs. The following are the default allowed usages, in case the GMs don’t say anything about it.

Phones may be used freely. Regular email may be used freely to communicate with individuals or groups of other Characters—not to mailing lists. Anonymous email and other tricks are not allowed (bcc is ok).

Zephyr may be used only to set up meetings with other Characters. Messages should be of the form “I’d like to meet with you and This-Person at This-Time at This-Place” and “Ok, see you there,” and may not convey any further information. In particular, you may not say what you want to discuss, as that often conveys important game information.

Finding Players with zlocate, finger, etc. is fine while you’re at the computer.

No other uses of computers are generally acceptable, including code-breaking or having the computer do anything for you in your absence.

When further Athena use is permitted, make sure you’re clear something is legal before doing it.

Remember the Athena Rules of Use; breaking into accounts is not permissible, and neither is looking at someone’s screen without their permission or other such privacy violations.

In some games you must cc all email to the GMs (generally gamename-gms@mit.edu). Hopefully the game rules will make it clear whether this is guaranteed to be secure or whether an ingame hacker might get to read your ingame email.

The GMs will use some list to send mail to the Players, generally rumor@mit.edu; Players may not send mail to this list without explicit GM approval (typically given for ingame newspapers publishing there).

’Kidnapping’ dropped; see 5.2 Bodies and 6.9.3 Captivity

Appendices

A Combat Systems While we include several combat systems here, there’s no need to read any but the one used by the game you’re playing. Always see 6.3 Martial Combat for generalities that hold for all these systems. None of these system descriptions say everything there is to know about martial combat in a specific game; the game rules will fill in the rest, such as an average person’s martial stats, the exact effects of weapons, etc. Remember that, as always, game-specific rules override this document.

A.1 MA / MD / HP This is the oldest of the combat systems described here. It involves three to four stats: Hit Points (HP; measures damage capacity), Martial Attack (MA; measures offensive ability), Martial Defense (MD; measures defensive ability), and possibly Martial Delta (Mdelta or MΔ; measures “flexible” ability). Whenever you are called upon to use your MA, MD, or MΔ, you may choose to “pull your punch” and use less than your full score (use integers, don’t go negative). Remember the basics are in 6.3 Martial Combat, including initiating, joining, escaping, and using weapons.

Tell your target what MA score you are using against them that round.

Add up all attacks being made against you that round; then subtract your Martial Defense. If the result is positive, you take that much damage; subtract it from your Hit Points. If the result is zero or less, congratulations; you defended so well you took no damage. Do not state your Martial Defense, your Hit Points, or how much damage you took, but do say whether you took any damage or not—the attacker(s) can tell if they got through.

If you were Surprised, your MA and MD are zero for the first round (only) of combat, regardless of any modifiers (Mdelta, weapons, …)—you get no attack and you take all the damage aimed at you.

If you reach 0 HP, you are knocked out (see 6.1 Health). When you awaken you’ll have 1 HP.

k.o.

When everyone has adjusted their Hit Points, the next round begins.

You may have a Martial Delta score, whether from special abilities, a weapon, or because everyone in the game does. If you have one, you use it to add to your MA and MD, splitting it between them as you wish at the start of each round, just before MA scores are declared. That is, if your Mdelta is 5, you can choose to take +5 MA, +0 MD, or +4 MA, +1 MD, or +3 MA, +2 MD, etc—any combination of plusses that adds to no more than 5. You may vary this from round to round.

The effects of weapons (see 6.3.1 Using Weapons) vary; they may give extra MA, MD, or MΔ, and may give skilled wielders even more. Typical unskilled benefits are Club +1 MA, Knife/Sword +2 MA.

Since by default martial stats cannot be stated, you cannot plan out group attacks numerically, though saying things like “let’s be as aggressive/defensive as possible” (if you have Mdelta) is legitimate.

Game rules will give typical martial stats, publicly known weapons effects, and healing rates.

Example Batman and Robin walk up to the Joker and declare “Martial Attack” together; he responds “Martial Defense” immediately. Batman’s stats are MA 12, MD 10, HP 15. Robin has MA 10, MD 8, HP 10; the Joker has MA 11, MD 13, HP 14, and happens to have a Knife out that gives him +2 MA.

In the first round, Batman tells the Joker “12” for his MA and Robin tells him “10”; the Joker chooses to attack Robin and tells him “13” (his MA of 11, +2 for the Knife). The Joker adds the attacks against him and gets 22, subtracts his MD of 13 to get 9, and since this is positive takes that much damage from his HP of 14, leaving him with 5. Robin subtracts his 8 MD from the 13 attack against him and loses 5 HP, leaving him with 5. Batman has not been attacked.

As everyone is still up and not trying to Break, the combat continues. Batman and Robin each attack the Joker again, and the Joker still attacks Robin. The attacks and defenses come out the same way, as no Mdelta was involved, so Robin takes 5 more HP, leaving him at 0, and the Joker takes 9, leaving him at -4. Both Robin and the Joker are unconscious; as Batman is the only combatant left the fight is over.

A.1 A.2 Rock, Paper, Scissors This system is based on the game “Rock, Paper, Scissors,” which involves showing hand

symbols to each other. It involves three stats: Hit Points (HP, measuring damage capacity), Martial Attack (MA, measuring offensive ability), and Martial Defense (MD, measuring defensive ability). You may always “pull your punch” by using less than your full MA or MD (integers only, no negatives). Remember the basics are in 6.3 Martial Combat, including initiating, joining, escaping, and using weapons in combats.

There are three normal symbols (the only ones in the real-world game):

Rock is made by a fist.

Paper is made by a flat hand held horizontally, palm down.

Scissors is made by the first two fingers extended out of a fist, looking vaguely like a pair of scissors. These represent normal, everyday combat maneuvers, and everyone can use them. They defeat each

other cyclically: Scissors cut Paper, Paper wraps Rock, Rock crushes Scissors. (Each ties itself.)

There are also three extra symbols, which represent more unusual things. You can’t use these unless you know you can. They don’t even always exist in every game using this system.

Cheat is a thumb’s-up gesture (thumb extended upward from fist). It defeats any of the normal three symbols, and ties another Cheat. Cheats represent things like martial arts, military combat training, unusual strength or speed, and so on. An ability to use Cheats will be of the form “2 Cheats per combat.” This means that between engaging in combat and disengaging, you have a total of 2 Cheats to use in all the rounds of combat; next time you get into a combat you again have 2 Cheats. If you don’t have such an ability, you can’t use Cheat.

Block is a flat hand held vertically, palm towards opponent. It represents a special defensive maneuver. Blocks defeat the three normal symbols and Cheats, and tie Blocks, but never do damage. Again, abilities to use these will be of the form “2 Blocks per combat.” If you don’t have such an ability, you can’t use Block.

Special is an “OK” gesture (index finger touching thumb in a circle, other three fingers extended, hand held vertically, palm outward). This represents something Weird, like magic. Every Special is different; if you have one you’ll know what it does, and if someone uses one on you they’ll tell you what happens. Unless the Special specifies otherwise, it ties any other symbol. There will usually be some additional effect like “You fall unconscious” or “You die.” Again, abilities to use these are usually of the form “2 Specials per combat.” If you don’t have such an ability, you can’t use Special.

Each combat round, count off “1, 2, 3, show.” On “show” each combatant displays a symbol towards the person they are attacking. You only attack one person, but you use your symbol to defend against everyone attacking you. If you were Surprised, you don’t show a symbol for the first round. Any time you don’t show a symbol (Surprised, or tied up, or…), you do no damage and lose to everyone attacking you.

If your symbol beats your target’s, you get to hurt them. Tell them your Martial Attack score. They subtract their Martial Defense score. If the result is positive, they lose that many Hit Points. You don’t get to know your target’s Martial Defense or Hit Points, but you can tell whether you hurt them at all or not.

Often games set everyone’s MA to 1 and MD to 0, so that a hit always does exactly 1 HP (+weapons). If multiple people attacked you, you use your symbol to defend against each separately (but only coun-

terattack one); for each attacker that beats your symbol, you take damage.

Blocks never do damage; they only protect you. (Their point is to take up your opponent’s limited number of Cheats and Specials and/or let you escape the combat.)

Unless otherwise specified, no one takes damage from a tie.

If you reach 0 HP, you are knocked out (see 6.1 Health). When you awaken you’ll have 1 HP.

k.o.

When everyone knows how much damage they took, the next round begins with “1, 2, 3, show” again. By default, symbols represent Character maneuvers and so you may coordinate attacks (see example). Game rules will give typical stats, publicly known effects of weapons and armor, and healing rates.

Example Batman and Robin both say “Martial Attack” to the Joker at once; he responds “Martial Defense.” They count “one, two, three, show” and display symbols: Batman shows a Rock towards the Joker and Robin a Paper (having agreed beforehand on that, so they’d be more likely to damage him); the Joker shows Paper towards Robin. The Joker and Robin tie, so neither of them damages the other. Batman’s Rock is defeated by the Joker’s Paper, so Batman does no damage; the Joker doesn’t damage Batman, since he wasn’t attacking him, even though Paper beats Rock.

Everyone counts “one, two, three, show” and redisplays. This round, Batman shows Paper and Robin Rocks, and the Joker switches to attacking Batman with a Scissors. Now the Joker damages Batman and Robin damages the Joker. The Joker says his MA score, “3”; Batman mentally subtracts his MD of 1, getting 2, and takes 2 HP. Robin says his MA score, “2”; the Joker subtracts his MD of 1, getting 1, and takes 1 HP. Everyone had 3 HP to begin with, so they’re all still up, and the fight continues.

A.2 A.3 Phi This evolved from the “Rock, Paper, Scissors” system. Remember the basics are in 6.3 Martial Combat,

including initiating, joining, escaping, and using weapons in combats.

You have one combat stat, Phi (the Greek letter ϕ). It serves simultaneously as offensive power, defensive power, and damage capacity. Memorize your Phi.

Rounds are based on the game “Rock, Paper, Scissors,” which involves showing hand symbols to each other. There are three symbols:

Rock is made by a fist.

Paper is made by a flat hand held horizontally, palm down.

Scissors is made by the first two fingers extended out of a fist, looking vaguely like a pair of scissors. These represent different kinds of combat maneuvers. They defeat each other cyclically: Scissors cut

Paper, Paper wraps Rock, Rock crushes Scissors. (Each ties itself.)

In each round of combat, count off “1, 2, 3, show.” On “show” each combatant displays a symbol towards the person they are attacking. You only attack one person, but you use your symbol to defend against everyone attacking you. If you were Surprised, you don’t show a symbol for the first round.

If you are wielding a weapon (see 6.3.1 Using Weapons), it adds its “weapon-like” Phi to yours. This will be on its item card; typical “weapon-like” Phi values are Club +1, Knife +2, Sword +3.

If your symbol beats your target’s, tell them your Phi. You may “pull your punch” and use a lower score. If an attacker wins RPS over a defender who has equal or higher Phi, the defender just loses one personal

(not weapon-like!) Phi.

If an attacker wins RPS over a defender who has lower Phi, the defender loses one personal Phi and is

k.o.

knocked out (see 6.1 Health).

On a tie in RPS, each defender loses one personal Phi (in a one-on-one fight, that’s both of you).

Remember that you use your symbol to defend against everyone, but you only attack the single target you extend your symbol towards.

If you don’t show a symbol (e.g. were Surprised), your Phi is zero (so you will be knocked unconscious). If your personal Phi reaches zero, you are knocked out (see 6.1 Health) immediately, weapon-like Phi

k.o.

notwithstanding. Phi never goes below zero.

After Phi scores are adjusted, a new round of combat begins with a new count of “one, two, three, show.” By default, symbols represent Character maneuvers and so you may coordinate attacks (see example). Phi lost in martial combat returns at a rate of one per minute (though not during combat). If you have

been knocked out in martial combat, you can only recover to Phi−1. (If it happens again, only to Phi−2… etc. If it happens so much you can’t recover beyond zero, you’re in a coma; tell a GM.) The remainder is long-term fatigue damage. It comes back at one Phi per night. Medical attention might also restore long-term damage; those with appropriate skills will know how this works.

Game rules will give typical personal Phi and publicly known weapon-like Phis.

Example Batman (personal Phi normally 4, but down to 2 from losing earlier combats), armed with a Bat-Knife (2 weapon-like Phi), and Robin (Phi 2, unarmed), both say “Martial Attack” to the Joker (Phi 3 and armed with a stolen Bat-Knife), who responds “Martial Defense.”

They count “one, two, three, show.” Batman extends a Rock towards the Joker and Robin a Scissors (having agreed beforehand on those, so they’d be more likely to damage him), while the Joker shows a Rock towards Robin. Batman and the Joker tie, so the Joker takes 1 Phi—Batman doesn’t, since the Joker wasn’t attacking him. The Joker’s Rock beats Robin’s Scissors, so the Joker tells Robin his personal+weapon Phi, “5”; Robin mentally compares this to his personal+weapon Phi of 2 and falls down unconscious. Batman is still at 2 personal + 2 weapon Phi; the Joker is now down to 2 personal + 2 weapon Phi also.

They again count “one, two, three, show.” Batman uses a Rock again and the Joker switches to Paper, winning. The Joker says his personal+weapon Phi, “4”; Batman mentally compares to his personal+weapon Phi of 4 and loses one personal Phi, bringing him down to 1.

In the third round, Batman switches to Paper, and the Joker stays with Paper. They tie, so each loses one personal Phi. The Joker is now at 1 personal Phi, so he’s still conscious. Batman is now at zero personal Phi, so even though he has 2 weapon Phi he falls unconscious. Woe to the Dynamic Duo! If the Joker leaves them alive, they’ll wake with 1 Phi each; since they’ve each lost a long-term Phi from fatigue, they each only recover to 1 Phi today anyway, so they’ll be pretty weak until they get some sleep or see a doctor. Meanwhile, the Joker gets his 2 lost personal Phi back at 1 per minute.

A.3 A.4 Drive-By Darkwater The original form of this system was written for the one-nighter Pirates of Dark Water.

new

The form presented here aims at maximum compatibility with observed variants of the system, i.e. it should be easy for GMs to effectively reproduce those by citing this and (where necessary) stipulating a few tweaks. Also see A.5 Melee Darkwater for a form with single-round instead of hit-and-run structure.

Most of the basic material in 6.3 Martial Combat holds, including combat initiation and use of weapons, but some things are overridden (especially defending against and helping with attacks) by the text below.

Your Combat Rating (CR) is your only combat stat. It serves as both offensive and defensive capability. When your CR is called for you may always “pull your punch” by using a lower value.

Attack When you declare an attack, append your CR: “Martial Attack 3!”

Defense If your rating is at least the attack’s, say “Resist” within two seconds and the attack fails.

If you haven’t said “Resist” two seconds after the attack (whether it beat you or surprised you), it succeeds. You are either knocked out or, if the attacker used a deadly weapon, wounded (see 6.1 Health). By default, deadly weapons are edged ones (the GMs may redefine this); these may have distinctive phsyreps.

Assist Allies within ZoC of both attacker and target may declare “Assist rating” at the attacker, who may then redeclare their attack with (up to) those added. (“Martial Attack 3,” “Assist 2,” “Martial Attack 5.”) If they do, the target then defends against the new attack instead of the original. All Assists must be within two seconds of the original attack declaration, and the redeclaration must be within a further two seconds.

Only the rating of the attack is changed, not of the attacker; they still rely on just their own CR to Resist any attacks directed at them. Mutually taking each other down is quite possible with Assists involved.

Assist is the only way to gang up. If several people each declare their own attack at someone, the attacks are resolved individually, not summed. The defender determines the order to resolve them if it’s unclear.

Special Attacks Some people may have abilities for special attack modes such as Restrain. These are used by replacing “Attack” with the alternate mode’s name (“Martial Restrain 3!”) when declaring any attack. Resolution proceeds normally, but the effects of success are replaced by those given by the special mode.

Possible alternate attacks include Restrain: You grab the defender, rendering them immobilized (see 6.1 Health) as long as you hold on. Disarm: While declaring, point to one item the defender holds. They drop it. Throw: While declaring, point in some direction. The defender is hurled ten full steps that way. Knockout: The defender is knocked out. Wound: The defender is wounded. There may be other modes, with or without public knowledge. Some modes may have prerequisites.

Special attacks can still be Assisted by anyone, whether or not they’re capable of the same attack mode. Only the rating changes when the Assisted attack is redeclared—the mode stays the same.

Drive-By Combat Each attack has its own existence and resolution, independent of its maker and all other attacks, including counter-attacks. You never become “engaged” as there is never a “round” of combat. Throwing a fist is like throwing a rock—once started on its way, it continues on its own without you.

Exactly one restriction is imposed on attackers: you may neither attack nor Assist for two seconds after doing either, except to redeclare the Assisted attack. You’re as free to fire a gun, run away, etc, as you would be otherwise, after or even while you declare. (Just be in ZoC during your declaration.)

Being attacked, however, restricts you to only martial action until you Resist it or two seconds pass. You cannot fire a gun, move around, nor do anything else save attack or Assist. Rapid serial attacks do not prevent other simple actions (drawing a weapon, firing, talking) in between.

Waylay If you have this ability (see 6.6 Waylay) you may combine it with any attack mode you have. Insert the mode into your announcement, e.g. “Martial Attack Waylay” or “Martial Restrain Waylay”; that mode’s effects then replace those of plain “Martial Waylay,” i.e. your victim is restrained (or whatever) immediately, without combat. You must still satisfy any prerequisites of the mode and your Waylay ability.

Example {T=0} Batman aims “Martial Attack 2” at the Joker. {T=1} The Joker responds “Martial Attack 3, Resist” and Robin calls “Assist 2,” all at Batman. {T=2} Batman declares “Martial Attack 4” at the Joker. {T=3} Batman falls; the Joker aims “Martial Attack 3” at Robin. {T=4} Joker falls. {T=5} Robin falls.

Robin should have run away at T=2, when Batman’s redeclaration bars the Joker from pursuing.

Had the Joker been armed, he could just Resist and then shoot Batman while Robin’s declaring Assist. Had the Dynamic Duo’s reaction times been much longer, Batman falls (at T=3) before starting to redeclare. Either way, once Batman goes down he can’t do anything, hence never gets to declare the Assisted attack.

A.4 A.5 Melee Darkwater The original form of this system was written for the one-nighter Pirates of Dark Water.

new

The form presented here aims at maximum compatibility with observed variants of the system (i.e. it should be easy for GMs to effectively reproduce those by citing this and (where necessary) stipulating a few tweaks) subject to having single-round structure in contrast to A.4 Drive-By Darkwater’s hit-and-run structure.

Remember the basics are in 6.3 Martial Combat, including initiation of and restrictions during combat. Your Combat Rating (CR) is your only combat stat. It serves as both offensive and defensive capability.

Attack When you declare an attack, append your CR: “Martial Attack 3!” You may “pull your punch” by using a lower value, but that one declared rating is both your attack and your defense in that combat.

Response The attacker’s target (and anyone else who wants to participate) has two seconds to respond with a Martial Attack of their own (against any combatant, not necessarily the original attacker/defender). Anyone failing to make a valid declaration in time has a zero rating for that combat.

Unless otherwise specified, 6.3.1 Using Weapons applies: you can respond with a ranged instead of a martial attack, but never both—if you fire you don’t declare, so any attackers not shot down will get you.

You cannot fire a gun that is being targeted by a Disarm attack (see below).

Resolution Everything resolves two seconds after the first declaration: for each combatant, if your rating is at least the sum of the attacks against you, they all fail; otherwise, they all succeed.

A successful attack leaves you either knocked out or, if with a deadly weapon, wounded (see 6.1 Health). By default, deadly weapons are edged ones (the GMs may redefine this); these may have distinctive phsyreps.

Note that with multiple weak attackers against a strong defender, it is quite possible for both the defender and the single attacker they counterattack to be taken down.

The combat is now over, even if several people remain standing. This means everyone may now draw a bigger weapon (or fire one) between combats, no matter how fast someone starts a new one.

Special Attacks Some people may have abilities for special attack modes such as Restrain. These are used by replacing “Attack” with the alternate mode’s name (“Martial Restrain 3!”) when making any declaration. Resolution proceeds normally, but the effects of success are replaced by those given by the special mode.

Possible alternate attacks include Restrain: You grab the defender, rendering them immobilized (see 6.1 Health) as long as you hold on. Disarm: While declaring, point to one item the defender holds. They drop it. Throw: While declaring, point in some direction. The defender is hurled ten full steps that way. Assist: (no effect; see below) Knockout: The defender is knocked out. Wound: The defender is wounded. There may be other modes, with or without public knowledge. Some modes may have prerequisites.

If attacks succeed whose effects contradict (Throw and Restrain, or Throws east and west) the defender is affected by whichever they think was declared earlier, or by whichever they choose if the order’s unclear.

Assist This special attack mode can be used by everyone; while its success has no effect, the attack does still add to the total against the defender, so by making it you help any other attacks against them.

Waylay If you have this ability (see 6.6 Waylay) you may combine it with any attack mode you have. Insert the mode into your announcement, e.g. “Martial Attack Waylay” or “Martial Restrain Waylay”; that mode’s effects then replace those of plain “Martial Waylay,” i.e. your victim is restrained (or whatever) immediately, without combat. You must still satisfy any prerequisites of the mode and your Waylay ability.

Example Batman tries gentle Bat Fu on the hatchet-wielding Joker, declaring “Martial Restrain 3.” Robin doesn’t know how to Restrain, so he chimes in with “Martial Assist 2” at the Joker instead. The Joker responds “Martial Attack 3,” pointing at Robin (having realized he’d just bounce off Batman’s equal rating).

The attacks against the Joker total 5, beating his 3, so they succeed. The attack against Robin is only 3, but that still beats his 2, hence it also succeeds. So Batman’s fine, and he holds the Joker immobilized, while Robin is wounded (since the hatchet is deadly; with a club he’d just be knocked out).

What if some henchman shoots Robin, say before the Joker targets him? Batman’s attack ties the Joker’s, so both fail. Since combat’s now over, Batman can draw his Bat-Knife even if the next starts immediately.

Had the Joker shot Robin, he’d still get Restrained, since then he couldn’t declare. Shooting Batman works better; now Robin’s attack succeeds, but that’s an Assist, harmless. If Robin uses plain Attack instead, the Joker needs to shoot both of them to escape intact.

A.5 A.6 Cinematic This system is slower but more entertaining to watch than the others; it is found in games that

want to discourage martial combat in favor of guns and/or that are of a genre involving lots of flashy fighting. Remember the basics are in 6.3 Martial Combat, including initiating, joining, and escaping combats. You only have one combat-relevant stat, your Hit Points, measuring damage capacity. Your offensive and

defensive capabilities are embodied in your Maneuver Cards. Each is a combat maneuver you can perform, whether you picked it up on the street, trained at the spy agency, or whatever.

Each round, each combatant selects one Maneuver Card and one other combatant. You use your chosen maneuver to defend against all attacks made against you that round as well as to attack your single target.

For each person attacking you (or that you’re attacking), hold up your card with the side that says “compare this side” facing you while they hold up theirs. Line up the cards. On the side of their card that you see there will be a line of boxes saying “Hit” or “Miss”; next to it, on your card, is a line of boxes saying “Block” or “No Block.”

Each “Hit” that lines up with a “No Block” does one Hit Point of damage to you. If the box says “Hit n” instead of “Hit,” you take n points of damage if your maneuver doesn’t Block it. (The attacker cannot “pull their punch” and punt Hits at this point.)

Of course, you only take damage from the people attacking you, and only do damage to the single person you are attacking, even in a large combat where you compare your maneuver to many other people’s.

You may coordinate maneuver choices with others, since they’re in-Character maneuvers and choices, but by default you may not discuss actual card box patterns (they’re like stats).

If you were Surprised, you get no maneuver the first round; you do no damage to anyone and take all the Hits on attackers’ cards.

If you are hit at all by an attacker’s maneuver, enact the effect described on the side of their card facing you. (Thus the “cinematic” nature of the system.) Some cards may have special effects beyond role-playing, like “you fall down” or “I escape.” Even if a maneuver that hits you has nothing special to role-play, your opponent can still tell if they hit you at all or not, though not how much they hit you.

Not all Maneuver Cards can be used at all times; they will list their requirements. For instance, a card that says “Requires: sword” cannot be used unless you are wielding a sword. (This is the point of weapons in this system; they let you use more studly maneuvers if you’re skilled with them.) Unless a card says otherwise, you must be standing to use it.

There are three maneuvers that anyone can use, even without a card. Pretend these maneuvers’ cards have all “Miss” and “No Block” on them; when you use one you do no damage to anyone and take all the Hits of anyone attacking you. These maneuvers are Stand Up (if you’re down), Fall Down (if you’re up) and Draw Weapon (no requirements). You declare that you are using such a maneuver at the same time as other maneuvers are revealed (by holding up cards); it takes effect at the end of the round (in case your opponent’s maneuver works only on fallen opponents, or whatever). Remember that you cannot fire a gun once engaged in martial combat.

If you reach 0 HP, you are knocked out (see 6.1 Health). When you awaken you’ll have 1 HP.

k.o.

After everyone has calculated their wounds and enacted any hit effects, the next round begins.

Game rules will give healing rates, and possibly typical numbers of Hits and Blocks.

Example

Batman sees

HIT HIT HIT miss miss

Patented Bat-Wham

A swirl of fists about the head and chest

Requires: nothing Compare this side {NO \atop BLOCK} block

{NO \atop BLOCK} block

{NO \atop BLOCK} HIT 3 miss miss miss miss Poke in the Eye with a Sharp Stick

funnier than a poke in the eye with a blunt stick

If hit: Clap hand over eye, exclaim “Ouch!”

Joker sees

HIT 3 miss miss miss miss Poke in the Eye with a Sharp Stick

funnier than a poke in the eye with a blunt stick

Requires: a sharp stick Compare this side

block

{NO \atop BLOCK} {NO \atop BLOCK} block block HIT HIT HIT miss miss

Patented Bat-Wham

A swirl of fists about the head and chest

If hit: Say “Oof!” and stagger back a step

Batman goes for a high-offense, moderate-defense Bat-Wham. Armed with a sharp stick, the Joker tries to poke Batman in the eye with it and defend with his other arm.

The Joker is hit twice, for one point each. Batman is hit only once, but for three points. The Joker says “Oof!” and staggers back, while Batman claps his hand over his eye and cries “Ouch!,” delighting onlookers.

If they both still have positive Hit Points, they now select their cards for the next round.

A.6 B Edition History This is a log of substantive changes made to the SR since the initial release. That almost always

means the same as “actual changes in the rules,” which only happen between editions (see Preface).

Each edition will internally flag the substantive changes between it and the previous edition by putting changebars in the margin, sometimes with associated marginal notes to clarify the nature of the change.

Non-substantive changes in wording or order, added cross-references, and changes to the L^{A}TEX source that don’t show up in hardcopy will often lack changebars and/or log entries.

1998–1999 Initial release. 1999–2000 No rules changes. 2000–2001

• Written Information has been rewritten for clarity to make several assumptions explicit.

• Martial Combat notes that maneuver choices may usually be coordinated. In the appendix, individual systems each note whether their choices are in-Character maneuvers and hence may be.

• Disarm and Cover was dropped.

• Running was dropped. 2001–2002

• Game Areas and Times dropped “soft” Game Breaks and Dead Times.

• Containers was added.

• Pickpocketing was dropped. 2002–2003 This edition never actually happened. 2003–2004

• 6.1 Health was added, defining knocked out and wounded, and those everywhere replace “become unconscious” and “bleed to death.” The associated waking/dying timers now default to five minutes. The term immobilized is also defined there.

• 6.9.3 Captivity was added (superseding part of Kidnapping, the rest becoming 5.2 Bodies).

• 6.4 Ranged Combat is completely rewritten. Limb shots are no longer distinguished by default; bleeding to death takes only five minutes; bleeding isn’t hastened by further hits.

• 6.5 Stealth Abilities and 6.6 Waylay were added.

• The Starlight combat system was dropped from the appendix.

• Combat systems A.4 Drive-By Darkwater and A.5 Melee Darkwater were added.

• 4.1 Restricted Areas was named, for the sake of games wanting to refer to that section specifically. ∘ There were lots of non-substantive changes, i.e. having no impact on what the actual rules are (hence

not requiring rereading) but hopefully making the document nicer. Various parts were reorganized (compare table of contents) and a few bits of text migrated among them. Many parts have been reworded for conciseness, clarity, emphasis, etc. The Preface’s material aimed at Assassins who predated the SR has been removed entirely, as has the general blurb on LARP.

2004–2005 No rules changes.

B.1